OpenStudy (anonymous):

What will be the magnetic force to a conducting spring that is near the infinite conducting wire. The wire attracts the spring but doesn't rotate it.

OpenStudy (anonymous):

HI.... \[\Huge \color{red}{\bigstar}~~\color{blue}{\ast~~\ast}~~\color{orange}{\bigstar}~~\color{blue}{\ast~~\ast}~~\color{purple}{\bigstar}~~\color{blue}{\ast~~\ast}~~\color{green}{\bigstar}~~\color{blue}{\ast~~\ast}~~\color{red}{\bigstar}\] \[\LARGE \color{red}{\mathbb{WELCOME}}~~\color{orange}{\mathbb{TO}}~~\color{purple}{\mathbb{OPENSTUDY}}~~\color{blue}{!!}~~~(:\] \[\Huge \color{red}{\bigstar}~~\color{blue}{\ast~~\ast}~~\color{orange}{\bigstar}~~\color{blue}{\ast~~\ast}~~\color{purple}{\bigstar}~~\color{blue}{\ast~~\ast}~~\color{green}{\bigstar}~~\color{blue}{\ast~~\ast}~~\color{red}{\bigstar}\]